OpenStudy (australia.j.):

Which of the following describes a linear function? a.) It is V shaped and it passes through the origin. b.) It is a straight line in one portion and a curve in another portion. c.) Its y-values increase at a non constant rate as its x-value increases. d.) Its y-values increase at a constant rate as its x-value decreases.

OpenStudy (australia.j.):

will medal and fan

OpenStudy (ikram002p):

well, linear function from its name its a line equation here a graph for three lines in three different colors 

OpenStudy (ikram002p):

a.) It is V shaped and it passes through the origin. well it need not to pass through the origin but it can through it sometimes. and V shaped is only for these type of equations with absolute value. |y=ax+b| b.) It is a straight line in one portion and a curve in another portion. a linear equation can never have a curve portion. c.) Its y-values increase at a non constant rate as its x-value increases. this give this equation y=axz+b and z is a variable rate which means its not linear equation , d.) Its y-values increase at a constant rate as its x-value decreases. this equation y=ax+b its linear enough =D

OpenStudy (australia.j.):

THANK YOU. @ikram002p

OpenStudy (ikram002p):

np :)